# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v2

@@ I.1 Additional Main Results
BC-PAR: BC-PAR [32] addresses dynamics mismatch through representation mismatch, measured as the deviation between the encoded source state-action pair and its next state. It employs a state encoder $f_{\psi}$ and a state-action encoder $g_{\xi}$, both trained on the target domain. The encoder loss is:
$\mathcal{L}(\psi,\xi)=\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\text{tar}}}\left[\left(g_{\xi}(f_{\psi}(s),a)-\text{SG}(f_{\psi}(s^{\prime}))\right)^{2}\right],$ (23)
where SG is the stop-gradient operator. Source rewards are adjusted as:
$\hat{r}_{\text{PAR}}=r_{\text{src}}-\beta\cdot\left\|g_{\xi}(f_{\psi}(s_{\text{src}}),a_{\text{src}})-f_{\psi}(s^{\prime}_{\text{src}})\right\|^{2},$ (24)
where $\beta$ controls the penalty strength. The actor ($\pi_{\varphi}$) and critic ($Q_{\varsigma_{i}}$) are jointly trained using both source and target data. Actor training includes a behavior cloning term, similar to BC-SAC. We implement BC-PAR using the official codebase, [^3] https://github.com/dmksjfl/PAR adopt the suggested hyperparameters, and train for 40K environment steps with 10 gradient updates per step.
[…]
@@
Optimizer Adam [24]
@@
Minimum log standard deviation $-20$
@@
Hyperparameter Source Flow Target Flow
@@
Table 4: Hyperparameter setup for source and target domain flows
@@ 18. Answer: [Yes]
  19. Justification: We provide all the experimental details in Appendix J.
@@ 20. Guidelines:
  23. Answer: [No]
  24. Justification: We will release the code upon acceptance. All datasets used in this paper are publicly accessible.
@@ 28. Answer: [Yes]
  29. Justification: We provide all the experimental details in Appendix J.
@@ 38. Answer: [Yes]
  39. Justification: All the computing infrastructure information has been provided Appendix J.
@@ 48. Answer: [Yes]
  49. Justification: We have discussed both potential positive societal impacts and negative societal impacts in Appendix A.

### v3

@@ Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
BC-PAR: BC-PAR [lyu2024cross] addresses dynamics mismatch through representation mismatch, measured as the deviation between the encoded source state-action pair and its next state. It employs a state encoder $f_{\psi}$ and a state-action encoder $g_{\xi}$, both trained on the target domain. The encoder loss is:
$\mathcal{L}(\psi,\xi)=\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\text{on}}}\left[\left(g_{\xi}(f_{\psi}(s),a)-\text{SG}(f_{\psi}(s^{\prime}))\right)^{2}\right],$ (44)
where SG is the stop-gradient operator. Rewards of the offline data are adjusted as:
$\hat{r}_{\text{PAR}}=r_{\text{off}}-\beta\cdot\left\|g_{\xi}(f_{\psi}(s_{\text{off}}),a_{\text{off}})-f_{\psi}(s^{\prime}_{\text{off}})\right\|^{2},$ (45)
where $\beta$ controls the penalty strength. The actor ($\pi_{\varphi}$) and critic ($Q_{\varsigma_{i}}$) are jointly trained using both offline and online data. Actor training includes a behavior cloning term, similar to BC-SAC. We implement BC-PAR using the official codebase, [^4] https://github.com/dmksjfl/PAR adopt the suggested hyperparameters, and train for 40K environment steps with 10 gradient updates per step.
CompFlow. When training the target flow, we use a quadratic cost function and employ the Python Optimal Transport (POT) library [flamary2021pot] to compute the optimal transport plan for each minibatch using the exact solver. Additional hyperparameters are provided in Table 3 and Table 4. Since the exploration bonus term is closely tied to properties of the environment—such as the state space, action space, and reward structure—it is expected that the optimal exploration strength $\beta$ varies across tasks. We perform a sweep over $\beta\in\{0.01,0.1,0.2\}$, and select the offline data selection ratio $\xi\%$ from $\{30\%,50\%\}$.
@@
Optimizer Adam [kingma2015adam]
@@
Minimum log standard deviation 20-20
@@
Hyperparameter Offline Flow Online Flow
@@
Table 4: Hyperparameter setup for the offline and online flows
Appendix L Experimental Details of Wildlife Conservation
We use the green security simulator in [xu2021robust]. The model is a Markov decision process with state, action, transitions, and a terminal return. We summarize the parts needed to reproduce our experiments.
State and action.
At time $t$, the state is
$s_{t}=\big(a_{t-1},\,w_{t-1},\,t\big),\qquad s_{0}=(0,\,w_{0},\,0),$
where $w_{t}=(w_{t}^{1},\dots,w_{t}^{N})$ is wildlife across $N$ cells and $a_{t}=(a_{t}^{1},\dots,a_{t}^{N})$ is patrol effort. The defender chooses $a_{t}\in[0,1]^{N}$ under the budget $\sum_{i=1}^{N}a_{t}^{i}\leq B$.
Adversary behavior.
At each step, the poacher places a snare in cell $i$ with probability
$p_{t}^{i}\;=\;\operatorname{logistic}\!\Big(z^{i}\;+\;\beta\,a_{t-1}^{i}\;+\;\eta\!\!\sum_{j\in\mathcal{N}(i)}a_{t-1}^{j}\Big),$
where $z^{i}$ is the baseline attractiveness of cell $i$. The parameters $\beta<0$ and $\eta>0$ capture deterrence from prior patrol and displacement from neighboring patrols. The realized attack is
$k_{t}^{i}\sim\mathrm{Bernoulli}\!\big(p_{t}^{i}\big).$
Wildlife transition.
After attacks, wildlife in each cell evolves by natural growth and poaching losses:
$w_{t}^{i}\;=\;\max\!\Big\{0,\;\big(w_{t-1}^{i}\big)^{\phi}\;-\;\alpha\,k_{t-1}^{i}\,\big(1-a_{t}^{i}\big)\Big\},$
where $\phi>1$ is the growth rate and $\alpha>0$ is the loss per uncovered attack. This defines the transition kernel $T_{z}$ over states given actions:
$s_{t+1}\sim T_{z}\!\big(s_{t},\,a_{t}\big).$
Return.
The episode return is the total wildlife at the horizon,
$R(s_{T})\;=\;\sum_{i=1}^{N}w_{T}^{i},$
and the expected return of a policy $\pi$ under environment parameters $z$ is
$r(\pi,z)\;=\;\mathbb{E}\!\left[R(s_{T})\right],\qquad s_{t+1}\sim T_{z}\!\big(s_{t},\,\pi(s_{t})\big),\;s_{0}=(0,w_{0},0).$
We assume access to an offline dataset of 100,000 transitions collected in Murchison Falls National Park using a well-trained SAC policy. For the online environment in Queen Elizabeth National Park, we are allowed a budget of 40,000 interactions.
Appendix M Discussion on Computational Cost
The practical cost of data filtering is relatively small. For example, with a batch size of $256$ for training the policy network and a Monte Carlo sample size of $30$, the entire filtering process takes just $0.03$ seconds on an A100 GPU. We will highlight in the paper that this efficiency is due to the simplicity of our flow training objective, which follows a linear path. As a result, solving the corresponding ODE at inference time is very easy—A basic Euler method with just $10$ time steps is sufficient.
@@ 18. Answer: [Yes]
  19. Justification: We provide all the experimental details in Appendix K.
@@ 20. Guidelines:
  23. Answer: [Yes]
  24. Justification: We have provided the code in the paper.
@@ 28. Answer: [Yes]
  29. Justification: We provide all the experimental details in Appendix K.
@@ 38. Answer: [Yes]
  39. Justification: All the computing infrastructure information has been provided Appendix K.
@@ 48. Answer: [Yes]
  49. Justification: We have discussed both potential positive societal impacts and negative societal impacts in Appendix B.